### 09 Apr 2026 v1

@@ 4.3 Method-level Post-Filtering
Semantic Alignment Score Computation: SAS quantifies semantic similarity between a method pair ($m_{1}$, $m_{2}$) from two projects. It compares class names, method names, API docstrings, and inline comments, and comprises three components: ① Class similarity ($simClass$) based on class names and docstrings; ② Method header similarity ($simMethodHeader$) based on method names, return type, parameter names and types; ③ Optional detail similarity ($simOptional$) based on names and types of local variable, method docstrings, and inline comments.
Each component is computed using the longest common subsequence (LCS) between token sequences, which has been shown effective for comparing code and documentation (Koznov et al., 2024; Roy and Cordy, 2008): $\text{sim}(S_{1},S_{2})=\frac{2\cdot\text{lcs}(S_{1},S_{2})}{n+m}$, where $S_{1}$ and $S_{2}$ have $n$ and $m$ tokens, respectively, and $lcs(S_{1},S_{2})$ is the length of their longest common subsequence. We compute LCS-based similarity for all relevant method details ($simClassName$, simClassDocsimClassDoc, simMethodNamesimMethodName, simReturnTypesimReturnType, simParamsimParam, simLocalVarsimLocalVar, simMethodDocsimMethodDoc, simCommentsimComment) and aggregate them into SAS components:
@@ 5.0.2 Evaluation Metrics
We evaluate baseline and our approach on dataset $D$ using the following metrics. In our context, true positives (TP) are genuine clones or code mappings in $D$ correctly detected, while false positives (FP) are non-clones or non-mappings incorrectly detected. True negatives (TN) are non-clones or non-mappings within $D$ correctly ignored, while false negatives ($FN$) are genuine clones or code mappings in $D$ missed by a detector. Method pairs outside $D$ are excluded from all calculations.
@@ 5.2.2 Post-filtering Effectiveness
Post-filtering on Raw Detection Results. To evaluate the practical impact of our post-filtering approach beyond labeled datasets, we applied a $SAS$ threshold of 0.5, the minimum optimal value observed in previous analysis, to the raw results detected by the selected techniques. Table 4 shows the filtered counts (“Filt”) and the percentage of removed pairs (“Out (%)”). We observed that DeepSim identified millions of production code clones, likely due to the large search space (e.g., comparing all pairs among 3,896 FindBugs and 3,970 SpotBugs methods in production code). The sheer volume makes manual verification impractical, highlighting the need for effective filtering. By filtering low-similarity clones, our rules help reduce the manual inspection cost of analyzing these clones. The average reduction rate is 33–99%.
@@ 7 Related Work
Code clone detection. Although some code clone detection tools target specific domains (e.g., smart contracts (He et al., 2020), apps (Wang et al., 2015), pre-training (Ding et al., 2023), or clone synchronization (Cheng et al., 2016)), we focuse on general-purpose code clone detection tools introduced in Section 2.2. In future, it is worthwhile to study enhancing domain-specific clone detection for redesigned projects. Prior approaches typically rely on abstract representations (e.g., program dependence graphs) or code normalization to capture similarity, but often ignore identifier names and code comments, which are crucial in redesigned projects. Documentation contains valuable domain knowledge, yet many tools ignore it due to the ambiguity of natural language. Several studies have confirmed its usefulness for code clone detection (Kuttal and Ghosh, 2020; Gupta and Goyal, 2021), e.g., Kuttal et al (Kuttal and Ghosh, 2020) use Latent Dirichlet Allocation (LDA) at file level combined with PDGs for function-level detection, while Gupta and Goyal (Gupta and Goyal, 2021) extract method documentation and apply Latent Semantic Indexing (LSI) with various similarity measures to identify concept-level clones. In contrast, our approach leverages near-duplicate documentation in redesigned projects using a simpler LCS-based method (Koznov et al., 2024), and combines it with identifier information. Moreover, our rules can be applied to prior code clone detection tools for further effectiveness enhancement.

### 10 Apr 2026 v2

@@ 4.3 Method-level Post-Filtering
Semantic Alignment Score Computation: SAS quantifies semantic similarity between a method pair ($m_{1}$, $m_{2}$) from two projects. It compares class names, method names, API docstrings, and inline comments, and comprises three components: ① Class similarity (simClasssimClass) based on class names and docstrings; ② Method header similarity (simMethodHeadersimMethodHeader) based on method names, return type, parameter names and types; ③ Optional detail similarity (simOptionalsimOptional) based on names and types of local variable, method docstrings, and inline comments.
Each component is computed using the longest common subsequence (LCS) between token sequences, which has been shown effective for comparing code and documentation (Koznov et al., 2024; Roy and Cordy, 2008): $\text{sim}(S_{1},S_{2})=\frac{2\cdot\text{lcs}(S_{1},S_{2})}{n+m}$, where $S_{1}$ and $S_{2}$ have $n$ and $m$ tokens, respectively, and $lcs(S_{1},S_{2})$ is the length of their longest common subsequence. We compute LCS-based similarity for all relevant method details (simClassNamesimClassName, simClassDocsimClassDoc, simMethodNamesimMethodName, simReturnTypesimReturnType, simParamsimParam, simLocalVarsimLocalVar, simMethodDocsimMethodDoc, simCommentsimComment) and aggregate them into SAS components:
@@ 5.0.2 Evaluation Metrics
We evaluate baseline and our approach on dataset $D$ using the following metrics. In our context, true positives (TP) are genuine clones or code mappings in $D$ correctly detected, while false positives (FP) are non-clones or non-mappings incorrectly detected. True negatives (TN) are non-clones or non-mappings within $D$ correctly ignored, while false negatives (FNFN) are genuine clones or code mappings in $D$ missed by a detector. Method pairs outside $D$ are excluded from all calculations.
@@ 5.2.2 Post-filtering Effectiveness
Post-filtering on Raw Detection Results. To evaluate the practical impact of our post-filtering approach beyond labeled datasets, we applied a SASSAS threshold of 0.5, the minimum optimal value observed in previous analysis, to the raw results detected by the selected techniques. Table 4 shows the filtered counts (“Filt”) and the percentage of removed pairs (“Out (%)”). We observed that DeepSim identified millions of production code clones, likely due to the large search space (e.g., comparing all pairs among 3,896 FindBugs and 3,970 SpotBugs methods in production code). The sheer volume makes manual verification impractical, highlighting the need for effective filtering. By filtering low-similarity clones, our rules help reduce the manual inspection cost of analyzing these clones. The average reduction rate is 33–99%.
@@ 7 Related Work
Code clone detection. Although some code clone detection tools target specific domains (e.g., smart contracts (He et al., 2020), apps (Wang et al., 2015), pre-training (Ding et al., 2023), or clone synchronization (Cheng et al., 2016)), we focuse on general-purpose code clone detection tools introduced in Section $2.$ In future, it is worthwhile to study enhancing domain-specific clone detection for redesigned projects. Prior approaches typically rely on abstract representations (e.g., program dependence graphs) or code normalization to capture similarity, but often ignore identifier names and code comments, which are crucial in redesigned projects. Documentation contains valuable domain knowledge, yet many tools ignore it due to the ambiguity of natural language. Several studies have confirmed its usefulness for code clone detection (Kuttal and Ghosh, 2020; Gupta and Goyal, 2021), e.g., Kuttal et al (Kuttal and Ghosh, 2020) use Latent Dirichlet Allocation (LDA) at file level combined with PDGs for function-level detection, while Gupta and Goyal (Gupta and Goyal, 2021) extract method documentation and apply Latent Semantic Indexing (LSI) with various similarity measures to identify concept-level clones. In contrast, our approach leverages near-duplicate documentation in redesigned projects using a simpler LCS-based method (Koznov et al., 2024), and combines it with identifier information. Moreover, our rules can be applied to prior code clone detection tools for further effectiveness enhancement.